Please help solve this eqaution:
1/4,-1/2 and 3/4,4
Find the equation of the line containing the given pair of points.
What is the equation of the line?
I am not very good when it comes to fractions.

Answer by scott8148(6628)

the y value changes 9/2 (from -1/2 to 4) ; while the x value changes 1/2 (from 1/4 to 3/4)

so the slope of the line is ___ (9/2) / (1/2) ___ or 9

y = 9x + b ___ plug in one of the points to find b

4 = 9(3/4) + b ___ 16/4 = 27/4 + b ___ -11/4 = b

y = 9x - 11/4